### v2

# Revealing Actual Viscoelastic Relaxation Times in Capillary Breakup

Abstract
We use experiments and theory to elucidate the size effect in capillary breakup rheometry, where pre-stretching in the visco-capillary stage causes the apparent relaxation time to be consistently smaller than the actual value. We propose a method accounting for both the experimental size and the finite extensibility of polymers to extract the actual relaxation time. A phase diagram characterizes the expected measurement variability and delineates scaling law conditions. The results refine capillary breakup rheometry for viscoelastic fluids and advance the understanding of breakup dynamics across scales.
preprint: APS/123-QED
Capillary thinning and breakup of viscoelastic liquid threads are ubiquitous in natural environments, biological tissues, and industrial applications [1, 2, 3]. Unlike viscous liquids, the capillary breakup of viscoelastic liquids exhibits distinct behaviors such as the exponential decay of filament thinning [4, 5, 6, 7] and the formation of satellite droplets on a string [8, 9, 10, 11], which are attributed to the elastic forces from polymeric stresses. Consequently, capillary breakup rheometry, which leverages these dynamic characteristics, has been a method widely used to measure viscoelastic properties, and is especially promising for weakly-elastic fluids [12] and biopolymers such as DNA [13, 14], whose properties are difficult to measure using common rheometric methods. Current well-developed experimental techniques include stretching a liquid bridge between two plates (known as CaBER) [15, 16, 17, 18], jetting [19, 20, 21, 22], dripping [23, 24], and dripping-onto-substrate (DoS) [25, 26, 27, 28].
Typically, the capillary breakup of viscoelastic fluids is governed by the interplay of inertial, viscous, elastic, and surface tension forces. Based on the sequence of events during the breakup of droplets or liquid bridges, the dynamics can be categorized into the visco-capillary (VC) or inertial-capillary (IC) stage, the elasto-capillary (EC) stage, and the final visco-elasto-capillary (VEC) stage [29]. By assuming infinite extensibility of the polymer chains ($b=\infty$) with a (longest) relaxation time $\lambda$, the evolution of the neck diameter $D(t)$ for the EC stage is well-known as
$D(t)/D_{1}=\left(GD_{1}/2\gamma\right)^{1/3}\exp\left[-\left(t-t_{1}\right)/3\lambda\right],$ (1)
where $G$ is the modulus of the viscoelastic fluid, $D_{1}$ is the initial diameter of the liquid filament, $\gamma$ is the surface tension, and $t_{1}$ is the time of onset of the EC stage [4]. This formula has been adopted in numerous studies to interpret the experimental results of the variation of the filament diameter $D(t)$ during the EC stage. Thus, one obtains an estimate of the apparent relaxation time $\lambda_{e}$ for various kinds of polymer solutions by fitting the exponential decay [15, 12, 30, 31, 25, 26, 27, 29, 32, 33].
Figure 1: Capillary breakup rheometry. (a) Schematic illustrations of liquid filament breakup for different experimental configurations: (i) dripping, (ii) dripping-onto-substrate (DoS), and (iii) a liquid bridge between two plates (classic CaBER). The liquid is initially held quasi-steady by a nozzle or plates with diameter $\ell$. The minimum neck diameter of the liquid, $D(t)$, evolves over time from the $D_{0}$ to $D_{1}$, where subscripts 0 and 1 mark the onset of the VC and EC stages, respectively. (b) Representative measurements for a PEO-PEG-8M solution using (i) dripping and (ii) DoS methods with a 13G nozzle ($\ell=2.4\,\mathrm{mm}$). (iii) and (iv) are the extracted evolutions of interface shapes with increasing time $t$ corresponding, respectively, to dripping and DoS, where $D_{0}$ is estimated by taking the filament aspect ratio $1\lesssim L/D_{0}\lesssim 3$. Scale bars represent 1 mm. (c) Representative datasets showing the variation of $D(t)$ for different $\ell\in\{0.3,0.9,2.4\}\,\mathrm{mm}$ and methods (different symbols). Distinct slopes of $\log_{10}D$ in the EC stage for varying $\ell$ indicate a size-dependent apparent relaxation time $\lambda_{e}$ predicted by the slope $\sim(3\lambda_{e})^{-1}$.
However, existing studies have found that pre-stretching of the polymers or filament thinning during the IC/VC stage significantly influences the exponential decay in the subsequent EC stage [14, 17, 18, 23, 32]. The sensitivity of $\lambda_{e}$ to these effects can be attributed to several experimental parameters, including the initial separation rate of the plates [17], the initial diameter $D_{0}$ of the liquid bridge [18] in the CaBER method, nozzle diameter $\ell$ in the dripping method (where $\lambda_{e}\propto\ell^{8/9}$ has been observed) [23], and the polymer extensibility parameter $b$ in the DoS method [32]. This parameter dependence introduces strong sensitivity of the extracted relaxation time $\lambda_{e}$ to initial conditions and inevitable pre-stretching effects in capillary breakup rheometry. While several studies have reported such dependence [23, 18, 32], no work has yet directly identified the actual relaxation time $\lambda$ from breakup measurements, nor established a general scaling law relating $\lambda_{e}$ to $\ell$ or $D_{0}$. In contrast, a recent study [33] reported no systematic dependence of $\lambda_{e}$ on CaBER plate size across four polymer solutions with varying concentration and molecular weight. This highlights the need for a unified framework to account for the size-dependent variability of $\lambda_{e}$ observed in different experimental systems.
In this Letter we demonstrate that size-dependent apparent relaxation times, $\lambda_{e}$, obtained from capillary breakup processes of varying scales, can be used to estimate the actual relaxation time, $\lambda$, while simultaneously determining the extensibility, $b$, representative of the polymer solution. In particular, accounting for finite extensibility $b$, we show that pre-stretching during the VC stage shortens the exponential decay in the elasto-capillary stage, leading to $\lambda_{e}<\lambda$, and that this effect can be dependent on the size of the experimental setups. Through experiments and numerical solutions, we unravel this quantitative relationship, enabling the determination of the actual relaxation time $\lambda$ from both our measurements and those reported in recent studies.
The capillary breakup processes considered here include dripping, DoS, and CaBER (Fig. 1a (i, ii, iii)). In all cases, the liquid initially is slowly stretched in a quasi-static manner within devices of characteristic size $\ell$. Under the action of capillary forces driving the Rayleigh-Plateau instability, the liquid rapidly contracts from the initial diameter $D_{0}$ of the VC stage to the initial diameter $D_{1}$ of the EC stage. In our experiments, we used needles with varying outer diameter $\ell\in\{0.3,0.5,0.9,1.8,2.4\}$ mm (corresponding to 30G, 25G, 20G, 15G, and 13G needles) to conduct tests using the dripping and DoS methods. Given that the influence of the size of the CaBER method has been tested in Ref. [18, 33], we also used their data of the PEOvis and PS solution (see [34], §I and Fig. S1) for our analysis.
Three kinds of polymer solutions were employed in our experiments: PIB-PB (polyisobutylene dissolved in a polybutene-mineral oil mixed solvent), PS-DOP (polystyrene dissolved in the organic solvent dioctyl phthalate), and PEO-PEG (polyethylene oxide dissolved in a polyethylene glycol-water solution), whereas PIB-PB-0.3 and PIB-PB-0.02 denote different PIB concentrations; PEO-PEG-8M and PEO-PEG-1M identify different molecular weights of PEO. These latter fluids can be regarded as Boger fluids with constant viscosity (Fig. S2 [34]), ensuring that additional effects from viscosity variations with extension rate are eliminated. Detailed preparation procedures and physical property measurements, such as viscosity $\eta_{0}$, surface tension $\gamma$, storage modulus $G^{\prime}$, and loss modulus $G^{\prime\prime}$, are provided in [34] §II and summarized in Table S1. The polymer concentrations $c$ used in this study fall within the dilute regime, with $c^{*}>c\gtrsim 0.1c^{*}$, and are all significantly higher than the minimal concentrations required for EC stage [35, 36]. These include $c_{\mathrm{min}}\sim c^{*}/b$ and $c_{\mathrm{low}}\sim c^{*}/b^{0.75}$, which represent the thresholds for the emergence of the EC regime [35] and for visibly observable exponential thinning [36], respectively.
Capillary breakup tests were conducted inside a transparent chamber to shield the system from air disturbances, with a liquid reservoir included to minimize the effects of evaporation (see [34] §III). High-speed imaging was conducted using a camera (VEO 640L, Phantom) equipped with microscope objectives (Nikon) of 2X, 5X, or 10X magnifications to accommodate varying nozzle diameters. A syringe pump (PHD 2000, Harvard Apparatus) controlled the flow rate at $Q=1-10\,\mu\mathrm{L/min}$, ensuring a quasi-static droplet for dripping. Taking the PEO-PEG-8M solution as an example, representative snapshots from the dripping and DoS tests, captured at a resolution of $2560\times 1600$ pixels and 1500 frames per second (fps), are shown in Fig. 1(b-i, ii), respectively. The liquid surface was analyzed using a custom MATLAB image-processing algorithm to track the gas-liquid interface as depicted in Fig. 1(b-iii, iv), in order to obtain the evolution of the minimum neck diameter $D(t)$ (Fig. 1c). Unlike the CaBER method, where the instability can be triggered in a controlled manner by adjusting the plate spacing to directly determine $D_{0}$, we propose that in the dripping and DoS methods, $D_{0}$ should be determined by ensuring $1\lesssim L/D_{0}\lesssim 3$ (Fig. 1b-iii, iv, detailed discussion seen in [34] §V). Each case was repeated at least three times to confirm data reproducibility. The results for the other solutions are provided in Fig. S4  [34].
Figure 2: Dependence of apparent relaxation time $\lambda_{e}$ on apparatus size $\ell$ for different Boger fluids using specific methods: PEO-PEG, PS-DOP, PIB-PB in this work by dripping and/or DoS, PEOvis (Ref. [18]) and PS16/PS7 (Ref. [33]) by CaBER.
The transition in the $D(t)$ curves and the appearance of a uniform filament (Fig. 1b-i, ii) allow the determination of the initial diameter $D_{1}$ for the EC stage. Using Eq. (1), a fit following this point (white lines in Fig. 1c) yields the value of $\lambda_{e}$. The results show that $\lambda_{e}$ increases with the nozzle diameter $\ell$, though no significant difference is observed between the $\lambda_{e}$ values obtained from the dripping and DoS methods. We conducted a comprehensive comparison of the apparent relaxation times $\lambda_{e}$ for different viscoelastic fluids, nozzle diameters $\ell$, and testing methods (Fig. 2). We observed that PEO-PEG, PIB-PB-0.02, and PEOvis solutions exhibit a size dependence of $\lambda_{e}$, showing a variation from $\sim\ell$ to $\sim\ell^{0.5}$ with increasing $\ell$ for one fluid. In contrast, $\lambda_{e}$ values of the PS-DOP and PIB-PB-0.3 solution remain nearly constant.
Figure 3: Prediction of the relaxation time ratio $\lambda_{e}/\mathrm{min}\{\lambda_{e}\}$ (a, c, e) and relative ratio $\lambda_{e}/\lambda$ (b, d, f) versus the initial liquid diameter $D_{0}$ for the PEO-PEG-8M solution, assuming $\lambda=$ 5 ms (a, b), 6 ms (c, d), and 9 ms (e, f), with varying extensibility $b\in\{2\times 10^{4},4\times 10^{4},8\times 10^{4}\}$. The symbols represent experimental results of $\lambda_{e}$ obtained from the dripping and DoS methods, suggesting a good approximation of $\lambda\approx 6$ ms and simultaneously indicating $b\approx 4\times 10^{4}$ for the PEO-PEG solution. (g) Comparison of experimental and numerical results using $(\lambda,b)=(6~{}\mathrm{ms},4\times 10^{4})$ for three different diameter needles.
We believe that viscoelastic fluids formed from nearly monodisperse polymer solutions should exhibit a constant intrinsic (longest) relaxation time $\lambda$, independent of testing conditions; indeed, this is the most common description of these solutions. To uncover the underlying dependence of $\lambda_{e}$ on experimental size $\ell$, we employed a zero-dimensional model based on the FENE-P model to predict the evolution of the dimensionless filament diameter $a(\tau)=D(\tau)/D_{0}$, according to equations for mass and momentum balances (subscripts $s$ and $p$ denote, respectively, solvent and polymer), and microstructure evolution along the axial ($z$) and radial ($r$) directions,
$\frac{Ec}{a}=\frac{3\beta}{1-\beta}E+f(A_{zz}-A_{rr})$ (2a)
$\partial_{\tau}A_{zz}=\left(2E-f\right)A_{zz}+1$ (2b)
$\partial_{\tau}A_{rr}=\left(-E-f\right)A_{rr}+1$ (2c)
$\partial_{\tau}a=-\frac{aE}{2},$ (2d)
where we have introduced dimensionless time $\tau=t/\lambda$, elasto-capillary number $Ec=2\lambda\gamma/\eta_{p}D_{0}$, solvent viscosity ratio $\beta=\eta_{s}/\eta_{0}$, dimensionless (time-varying) extensional strain rate $E=-2\lambda D^{-1}\frac{dD}{dt}$, the conformation tensor $\bm{A}$ with $rr$ and $zz$ components, and the finite extensibility function $f=(b-3)/(b-A_{zz}-2A_{rr})$. By comparing simplified analytical solutions (see [34] §IV-B1, B2), approximate solutions ([34] §IV-B3), and an analysis of the full model ([34] §IV-C1) under different assumptions, we demonstrated that the evolution is sensitive to the initial conformation $\bm{A}(t_{1})$ for the EC stage. Only by solving the full model numerically can the variations in the VC, EC, and VC-EC transitions be captured accurately (see [34] §IV).
With the initial conditions $a=A_{rr}=A_{zz}=1$, Eqs. (2d) can be solved numerically ([34] §IV-C2) using MATLAB’s ODE15s solver to yield the temporal evolution of the dimensionless diameter $a(\tau)$, conformation $A_{zz}$ and $A_{rr}$, and extensional strain rate $E(\tau)$ (typical results are shown in Fig. 3 for $a(\tau)$; see also Fig. S10 for all variables). Based on the numerical results, we calculate the numerically predicted $\lambda_{e}/\lambda=\max\left\{{2}/{(3E(t))}\right\}$ for $t\geq t_{1}$, which corresponds to the early period of the EC stage (see [34], Fig. S10d), under various parameter choices of $\beta$, $Ec$, and $b$. This further allows the construction of a phase diagram of $\lambda_{e}/\lambda$ as a function of the three parameters (see [34], Fig. S10e). Notably, every fixed $\beta$ enables the phase diagram to be approximated with a two-dimensional representation in terms of $Ec$ and $b$ (see [34], Fig. S10f).
For a specific viscoelastic liquid, as properties $\eta_{p}$, $\beta$ and $\gamma$ are measurable and size $D_{0}$ is controlled, we can assume a true relaxation time $\lambda$ for determining $Ec$ in Eq. (2a), and explore a wide range of $b$ values in the numerical calculations. Taking the PEO-PEG solution as an example, numerical predictions of the relative ratio $\lambda_{e}/\mathrm{min}\{\lambda_{e}\}$ (Fig. 3a, c, e) and $\lambda_{e}/\lambda$ (Fig. 3b, d, f) for assumed $\lambda\in\{5,6,7\}\,\mathrm{ms}$ and $b\in\{10^{4},2\times 10^{4},4\times 10^{4},8\times 10^{4}\}$ are compared with experimental results (symbols in Fig. 3). The trend of $\lambda_{e}/\mathrm{min}\{\lambda_{e}\}$ for different $D_{0}$ values provides a significant indicator for comparison with model predictions because both $\lambda_{e}/\lambda$ and $\lambda_{e}/\mathrm{min}\{\lambda_{e}\}$ are uniquely determined for fixed $Ec$ and $\beta$.
The comparisons of the two trends indicate that when $\lambda\approx 6\,\mathrm{ms}$, the numerical predictions align well with the experimental results (dash-dot curves in Fig. 3g). Furthermore, at the same time this analysis determines $b\approx 4\times 10^{4}$. We emphasize that with this approach, only $\lambda$ needs to be adjusted for comparison, while $b$ is determined through the matching of the two trends. The comparison of experimental and numerical $a(\tau)$ shown in Fig. 3g suggests a good estimation for $\lambda$ and $b$ has been achieved. The same procedure was used to find a good agreement with the experimental data from [18] (Fig. S11 in [34]).
Figure 4: Dimensionless map of the relaxation time ratio $\lambda_{e}/\lambda$ versus the measurable elasto-capillary number $Ec_{e}=2\gamma\lambda_{e}/(D_{0}\eta_{p})$ for various unknown extensibility values $b\in[10^{2},10^{6}]$. (a) 3D maps for different $\beta$ of a specific solution. (b) Top view of 3D maps. (c) Collapsed 2D map for easy comparison. The symbols represent experimental results (red: PEO-PEG-8M, purple: PEO-PEG-1M, green: PIB-PB-0.3, black: PIB-PB-0.02, blue: PS-DOP) from current work, Ref. [18] (gray: PEOvis), and Ref. [33] (light/dark orange: PS16-50/100, light/dark brown: PS7-200/500). Solid lines are obtained from fitting, with deviation ranges enclosed by dash-dot lines, suggesting the estimations.
Next, we analyzed the experimental data presented in Fig. 2 and constructed a dimensionless phase diagram (note the logarithmic axes) of $\lambda_{e}/\lambda$ as a function of $Ec_{e}$, $b$ and $\beta$ (Fig. 4a), where $Ec_{e}=2\gamma\lambda_{e}/(D_{0}\eta_{p})$, with each variable being directly measurable from experiments. Although different $\beta$ values affect the map, particularly for small $b$, indicated in Fig. 4b, we can construct the phase diagram with all experimental data as shown in Fig. 4c; uncertainties in all measurable parameters were considered to determine the true relaxation time $\lambda$ and the range of extensibility $b$ for each viscoelastic fluid tested.
For the PEO-PEG-8M solution, the relaxation time is estimated as $\lambda\approx 6\,\mathrm{ms}$, with $b=4\times 10^{4\pm 0.2}$. For the PEOvis solution, $\lambda\approx 150\,\mathrm{ms}$ and $b=2\times 10^{4\pm 0.1}$, consistent with the results from Ref. [18], where $b\approx 2\times 10^{4}$ was estimated using an additional model. This agreement reinforces the validity of our approach. Furthermore, the larger $b$ for PEO-PEG-8M compared to PEOvis aligns with their respective molecular weights, $M_{w}\sim 8\times 10^{6}$ in our experiments versus $M_{w}\sim 4\times 10^{6}$ in Ref. [18], given similar solvent mixtures (see [34], Tab. S1). This observation is further supported by the PEO-PEG-1M solution with $M_{w}\sim 1\times 10^{6}$, which yields $(\lambda,b)\approx(2.2~{}\mathrm{ms},7\times 10^{2\pm 0.2})$.
For the PIB-PB-0.3 solution, experimental uncertainties prevent a precise determination of whether $\lambda_{e}$ varies with $D_{0}$, suggesting that a reasonable estimation of $\lambda$ and $b$ is not readily attainable. However, utilizing the more dilute PIB-PB-0.02 solution, we estimate $(\lambda,b)\approx(700\,\mathrm{ms},2\times 10^{3\pm 0.1})$. Given that both solutions contain the same polymer at different dilutions, it is reasonable to assume that $b$ remains unchanged. This leads to an inferred relaxation time of $\lambda\approx 850\,\mathrm{ms}$ for PIB-PB-0.3.
From the phase diagram (Fig. 4), two distinct regions can be identified. When $\log_{10}(Ec_{e}^{-1}b)\gtrsim 1.5$, the experimentally measured $\lambda_{e}$ becomes independent of the experimental size, i.e. $\lambda_{e}\sim\ell^{0}$. However, this does not imply that the true relaxation time $\lambda$ is obtained. In this region, since $b$ cannot be experimentally fitted, an additional evaluation of $b$ is necessary to determine the specific value of $\lambda_{e}/\lambda$. In the region where $\log_{10}(\lambda_{e}/\lambda)\lesssim-0.6$, the slope of $\log_{10}(\lambda_{e}/\lambda)$ approaches infinity, indicating that $\lambda_{e}\sim D_{0}$. This is consistent with the scaling law $\lambda_{e}\sim\ell$ observed in Fig. 2 when $\ell(D_{0})$ becomes sufficiently small. The $\lambda_{e}\sim\ell^{0.5}$ scaling law observed in Fig. 2 represents an empirical scaling behavior characterizing the transition between the two regions.
Thus, some predictions can be made. For fluids with relatively low surface tension and strong viscoelasticity, such as PIB-PB-0.3 and PS-DOP, which exhibit $\log_{10}(Ec_{e})\lesssim 1.5$ during capillary breakup, a size-independent exponential decay in the EC stage may occur. This implies that a nearly constant apparent relaxation time can be observed. In contrast, fluids with relatively high surface tension and low viscoelasticity, such as water-based dilute solutions like PEO-PEG and PEOvis, exhibit significant size dependence of the apparent relaxation time. The extensibility $b$ determines the value of the critical $Ec_{e}$ at which this effect occurs, with smaller $b$ making this influence more readily triggered. Additionally, we speculate that in micro-scale capillary breakup processes (e.g., within biological tissues), where $Ec_{e}$ is extremely big, the linear dependence of $\lambda_{e}\sim D_{0}\sim\ell$ is likely to occur.
In summary, our study reveals, through experiments and numerical solutions, how inevitable pre-stretching during the visco-capillary stage of capillary breakup influences the exponential decay in the elasto-capillary stage and the apparent relaxation time $\lambda_{e}$ obtained using classical formulations for viscoelastic fluids. Through a theoretical model, we demonstrate that the apparent relaxation time is always smaller than the actual relaxation time, i.e., $\lambda_{e}<\lambda$. We further identify that the size dependence of $\lambda_{e}$ is governed by the extensibility $b$, solvent viscosity ratio $\beta$, and the dimensionless number $Ec_{e}=2\gamma\lambda_{e}/(D_{0}\eta_{p})$. Large $b$ and small $Ec_{e}$ result in size-independent behavior, while under big $Ec_{e}$ conditions, the scaling $\lambda_{e}\sim D_{0}\sim\ell$ emerges. We present a methodology for determining the actual relaxation time $\lambda$ by synthesizing a series of apparent relaxation times $\lambda_{e}$ measured under various size conditions. These findings provide a new perspective on accurately measuring viscoelastic properties using capillary breakup rheometry and offer insights into the size effects in the capillary breakup dynamics of viscoelastic fluids. Looking forward, several factors need to be considered to address the measurement of more complex fluids, such as those with shear-thinning viscosity, fluids where the IC stage dominates instead of the VC stage.
Acknowledgements.
We thank Evgeniy Boyko for valuable discussions. We thank the NSF-supported project (NSF-BSF: Explaining the Mismatch of Experiments and Simulations for Viscoelastic Flows, 2246791) and the Princeton Materials Research Science and Engineering Center (MRSEC, DMR-2011750) for partially funding this work. J.H. acknowledges the Kwanjeong Educational Foundation Graduate Fellowship for financial support.